OpenStudy (trojanpoem):

Find the work done on the ideal gas:

OpenStudy (trojanpoem):

Relation:

OpenStudy (trojanpoem):

Work done = area under the curve from A to B

Parth (parthkohli):

Yes, correct.

OpenStudy (trojanpoem):

Area of the segment + area of rectangle ?

Parth (parthkohli):

correct

OpenStudy (trojanpoem):

Now area of segment ? I remember that area of segment = Area of sector - area of triangle. but not quite sure how to apply this here.

Parth (parthkohli):

ah, I see. so you're not necessarily sure if it is a semicircle they've made.

OpenStudy (trojanpoem):

Not semi-circle as Radius y-axis = (500-300) * 10^3 = 2*10^5 Pa Radius x-axis = (6-3.6) * 10^-3 = 2.4 * 10^-3 m^3

Parth (parthkohli):

I know what you're saying, but they're completely different units anyway.

OpenStudy (trojanpoem):

and I doubt we can turn Pa into m^3

Parth (parthkohli):

to get a dimensionally correct answer\[\frac{1}2\pi \left(\frac{V_2 - V_1}2\right)\left(\frac{P_2 - P_1}2\right) \]apply the above

OpenStudy (trojanpoem):

V2 = 6 , V1 = 3.6 or V1 = 1.2

Parth (parthkohli):

If you see what I did, I just wrote the radius in two different ways as the formula is \(1/2 \pi r \times r\)

OpenStudy (trojanpoem):

r = (V2-V1)/2 = Diameter/2

OpenStudy (trojanpoem):

But how did you induce that ?

Parth (parthkohli):

Just to get a dimensionally correct answer. That's all. Try getting an answer this way.

OpenStudy (trojanpoem):

2193.6 correct.

Parth (parthkohli):

Oh did you get that answer?

OpenStudy (trojanpoem):

Used your formula and compared it to the book final answer which is 2194

Parth (parthkohli):

Good :)

OpenStudy (trojanpoem):

Thanks !